OpenStudy (anonymous):

Solve 4x = 12 The logarithmic form is log _4 12=x Where do I go from there?

OpenStudy (anonymous):

\[4^x=12\iff x=\frac{\ln(12)}{\ln(4)}\]

OpenStudy (anonymous):

Ok lemme put those in

OpenStudy (anonymous):

I got 2.15 but these are the choices? x ≈ 0.27 x ≈ 0.56 x ≈ 1.79 x ≈ 19.93